ROTATIONAL SPECTRUM OF CYANOPHOSPHAACETYLENE, H_{2}P-C≡C-C≡N

LU KANG, Department of Natural Sciences, Union College, Barbourville, KY 40906; and \underline{STEWART E. NOVICK}, Department of Chemistry, Wesleyan University, Middletown, CT 06459.

The a type transitions of the microwave rotational spectra of cyanophosphaacetylene, H_{2}P-C≡C-C≡N, have been investigated in the frequency region between 5 and 26.5 GHz by Fourier Transformation Microwave (FTMW) spectroscopy. Rotational, centrifugal distortion and ^{14}N nuclear quadrupole coupling constants have been determined. Density functional theory level ab initio calculations were performed to predict the molecular constants, and the predicted values are in good agreement with our experimentally determined results. The ^{13}C and ^{15}N isotopomer transitions were also observed. The derived r_{0} structure is quite comparable to the calculated H_{2}P-C≡C-C≡N equilibrium geometry.